OpenStudy (anonymous):

3 2 - 1--- (-3 --- ) 7 x 3 A. 2 3 -- 7 B. 2 -5 --- 21 C. 11 D. 5 5 --- 21

OpenStudy (kropot72):

First convert the mixed numbers to fractions:\[-1\frac{3}{7}=-\frac{10}{7}\] \[-3\frac{2}{3}=-\frac{11}{3}\] Now multiply the fractions: \[-\frac{10}{7}\times -\frac{11}{3}=?\]

OpenStudy (anonymous):

i got -10001331

OpenStudy (kropot72):

When the two negative signs are multiplied you get a positive sign for the answer. Now just multiply the numerators, 10 * 11= something, and multiply the denominators, 7 * 3 = something else. Now divide 'something' by 'something else' to get the answer.

OpenStudy (anonymous):

i got 5.2380952381 ?

OpenStudy (kropot72):

You are right. However the answer choices are either mixed numbers or an integer. If you divide 110 by 21 you need to do a manual long division calculation to get a mixed number answer (unlees your calculator handles fractions and mixed numbers). Can you try 110/21 using long division?

OpenStudy (anonymous):

idk how ? i only got my phone but i think the answer is 5 5 over 21

OpenStudy (kropot72):

Yes, good work! The correct choice is D: \[5\frac{5}{21} \]

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're most welcome :)